OpenStudy (anonymous):

2. Which number is prime? (Points : 4) 91 53 27 14

OpenStudy (anonymous):

Do you know the definition for prime numbers?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, so which number can you eliminate immediately?

OpenStudy (anonymous):

is the answer 14

OpenStudy (anonymous):

Is 14 divisible by anything *other* than 1 and itself?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I should say, "evenly divisible" (meaning when you divide, you get a whole number with no remainder) So yes, 14 is evenly divisible by 2 and also by 7 (14/2 = 7, 14/7 = 2)

OpenStudy (anonymous):

So we can eliminate 14 as a possible answer.

OpenStudy (anonymous):

A quick tip, any even number (a number ending in 0, 2, 4, 6, or 8) is not a prime number, because it is divisible by 2.

OpenStudy (anonymous):

So, let's look at 27 next. Do you know of anything that 27 is evenly divided by?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Exactly. 27/3 = 9

OpenStudy (anonymous):

So 27 is not a prime number.

OpenStudy (anonymous):

What about 53 and 91?

OpenStudy (freckles):

2 is actually the only even prime

OpenStudy (anonymous):

That's true, 2 is evenly divisible by itself and 1, but by nothing else, making it the only even prime number.

OpenStudy (anonymous):

So what about 53 .... is 53 evenly divisible by anything?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

You are correct, 53 is the prime number.

OpenStudy (anonymous):

91 is divisible by 7, 91/7 = 13.

OpenStudy (anonymous):

For future reference, this webpage may help you: http://www.mathgoodies.com/lessons/vol3/divisibility.html